a sample of iron reated with hydrochloric acid. the liberated hydrogen occupied 40.1ml when collected over water at 27 degree celsus at a barometric pressure of 750 torr. what is the mass of the sample of iron? ( Because of the collection of gas over water, the vapor pressure of the water must be subtrated from the barometric pressure to find the pressure of the hydrogen gas)
Answer #1

at 27oC,vapour pressure of water = 24 torr

source: https://en.wikipedia.org/wiki/Vapour_pressure_of_water

P (H2) = barometric pressure - vapour pressure of water = 750 - 24 = 726 torr

            = 726/760 atm

            =0.9553 atm

V = 40.1 mL = 0.0401 L

T= 27 oC = (273+27) K= 300 K

We need to calculate number of moles of H2 liberated,

use:

P*V=n*R*T

0.9553 * 0.0401 = n*0.0821*300

n=0.00156 mol

Reaction occuring here is:

Fe + 2HCl --> FeCl2 + H2

1 mol of H2 is liberated from 1 mol of Fe

so,

number of moles of Fe reacting = 0.00156 mol

molar mass of Fe = 56 gm

mass of Fe = molar mass * number of moles

                    = 56 * 0.00156

                    = 0.087 gm

Answer: 0.087 gm